OpenStudy (anonymous):

Write the given expression as an algebraic expression tan(2 sin^−1 x) in x

OpenStudy (anonymous):

look up the double angle formula for tangent i forget what it is but that is what you need

OpenStudy (anonymous):

tan(2arcsin(x)) To make it easier to see for myself.

OpenStudy (anonymous):

maybe \[\tan(2x)=\frac{2\tan(x)}{1-\tan^2(x)}\]? something like that?

OpenStudy (anonymous):

Yeah thats it.

OpenStudy (anonymous):

then \[\frac{2\tan(\arcsin(x)}{1-\tan^2(\arcsin(x)}\] is a start

OpenStudy (anonymous):

all you need is \(\tan(\arcsin(x))\) to finish

OpenStudy (anonymous):

Let y = arcsin(x) Then sin(y) = x, and tan(arc sin x) = tan(y)

OpenStudy (anonymous):

arcsin(sin(y)) = y

OpenStudy (anonymous):

\[= \frac{ 2\tan(y) }{ 1-\tan ^{2}(y) }\]

OpenStudy (anonymous):

this is what i get, is rite? tan(y) =[2x((√1-x^2)+x)]/(1-2x^2)

OpenStudy (anonymous):

I got \[\frac{ x }{ \sqrt{1-x ^{2}} }\]

OpenStudy (anonymous):

I've put your answer into my online assignment and it stated to be incorrect?

OpenStudy (anonymous):

Oh I've got it, it's [2x \sqrt{1-x^2}/(1-2x^2)\]

OpenStudy (anonymous):

\[(2x \sqrt{1-x^2})/(1-2x^2)\]